# Proton Dynamics Scenarios in the Integrable Optics Test Accelerator (IOTA) at Fermilab

### v1

@@ 7 Discussion and Conclusions
We can develop measures to mitigate the driving mechanisms of emittance growth and beam loss for various experiments. Baking the IOTA vacuum chamber above $100\,^{\circ}$C can reduce the residual gas content seen in Fig. 5, especially the partial pressure of water, leading to a better beam lifetime and reduction in emittance growth rates at low intensities. At intermediate intensities, incorporating an electron cooler[26] as a part of the electron lens program can compensate for IBS and lead to better control of the phase-space of the beam. At high-intensities where space-charge plays an important role, optimization of the bare lattice tune, inclusion of periodicity and compensation of Resonance Driving Terms[52, 53] can reduce emittance growth. For instance, compensating for the 4th order resonance in the DN lattice by turning on octupoles can greatly reduce resonance-driven diffusion at tune shifts $|\Delta\nu_{y,\mathrm{SC}}|\lesssim 0.3$ while keeping the linear optics unchanged. Beside conventional measures, the IOTA research program will also explore NIO to improve stability at resonances, and space-charge compensation using an electron lens[14, 15, 54] to reduce the effective tune footprint of the beam with broad applicability in high-intensity synchrotrons and storage rings.

### v2

@@ 7 Discussion and Conclusions
We can develop measures to mitigate the driving mechanisms of emittance growth and beam loss for various experiments. Baking the IOTA vacuum chamber above $100\,^{\circ}$C can reduce the residual gas content seen in Fig. 5, especially the partial pressure of water, leading to a better beam lifetime and reduction in emittance growth rates at low intensities. At intermediate intensities, incorporating an electron cooler [28] as a part of the electron lens program can compensate for IBS and lead to better control of the phase-space of the beam. At high-intensities where space-charge plays an important role, optimization of the bare tune (linear optics without space-charge), inclusion of periodicity and compensation of Resonance Driving Terms [54, 55] can reduce emittance growth. For instance, compensating for the 4th order resonance in the DN lattice by turning on octupoles can greatly reduce resonance-driven diffusion at tune shifts $|\Delta\nu_{y,\mathrm{SC}}|\lesssim 0.3$ while keeping the linear optics unchanged. Beside conventional measures, the IOTA research program will also explore NIO to improve stability at resonances, and space-charge compensation using an electron lens [14, 15, 56] to reduce the effective tune footprint of the beam with broad applicability in high-intensity synchrotrons and storage rings.